OpenStudy (anonymous):

Given that f(x) = 2x + 1 and g(x) = −5x + 2, solve for f(g(x)) when x = 3.

OpenStudy (anonymous):

f(g(x)) = 2(-5x +2) +1 = -10x + 4 +1 = -10x + 5 at x =5 => f(g(x)) = -10(3) + 5 = -30 +5 = -25

OpenStudy (anonymous):

sorry at x = 3=> f(g(x)) = -25